# Benchmarking Post-Training Quantization in LLMs: Comprehensive Taxonomy, Unified Evaluation, and Comparative Analysis

### v1

@@ Abstract
Post-training Quantization (PTQ) technique has been extensively adopted for large language models (LLMs) compression owing to its efficiency and low resource requirement. However, current research lacks a in-depth analysis of the superior and applicable scenarios of each PTQ strategy. In addition, existing algorithms focus primarily on performance, overlooking the trade-off among model size, performance, and quantization bitwidth. To mitigate these confusions, we provide a novel benchmark for LLMs PTQ in this paper. Firstly, in order to support our benchmark, we propose a comprehensive taxonomy for existing mainstream methods by scrutinizing their computational strategies (e.g., optimization-based, compensation-based, etc.) Then, we conduct extensive experiments with the baseline within each class, covering models with various sizes (7B-70B), bitwidths, training levels (LLaMA1/2/3/3.1), architectures (Mixtral, DeepSeekMoE and Mamba) and modality (LLaVA1.5 and VILA1.5) on a wide range of evaluation metrics.Through comparative analysis on the results, we summarize the superior of each PTQ strategy and modelsize-bitwidth trade-off considering the performance. For example, our benchmark reveals that compensation-based technique demonstrates outstanding cross-architecture robustness and extremely low-bit PTQ for ultra large models should be reexamined. Finally, we further accordingly claim that a practical combination of compensation and other PTQ strategy can achieve SOTA various robustness. We believe that our benchmark will provide valuable recommendations for the deployment of LLMs and future research on PTQ approaches.
Benchmarking Post-Training Quantization in LLMs: Comprehensive Taxonomy, Unified Evaluation, and Comparative Analysis
Jiaqi Zhao1thanks: Equal Contribution, Ming Wang [^1] footnotemark: 1, Miao Zhang1thanks: Corresponding Author, Yuzhang Shang [^2] footnotemark: 2, Xuebo Liu1, Yaowei Wang1, Min Zhang1, Liqiang Nie1 1 Harbin Institute of Technology (Shenzhen) 2 Illinois Institute of Technology jiaqizhao0455@outlook.com, 190110509@stu.hit.edu.cn, yshang4@hawk.iit.edu {zhangmiao, liuxuebo, wangyaowei, zhangmin2021}@hit.edu.cn, nieliqiang@gmail.com
@@ 1 Introduction
Large language models (LLMs) have achieved remarkable success in text generation and various reasoning tasks, with representative ChatGPT (Achiam et al., 2023) and LLaMA family (Touvron et al., 2023a, b; Dubey et al., 2024). However, their massive parameter scale imposes significant memory and inference overhead, which constrain their practical deployment. To address the issue, numerous model compression techniques have been proposed, such as quantization (Lee et al., 2024; Shang et al., 2024), pruning (Frantar and Alistarh, 2023; Cheng et al., 2024a, b), low-rank decomposition (Hu et al., 2021; Yuan et al., 2023), and knowledge-distillation (Gou et al., 2021). Among these, Post-training Quantization (PTQ) (Yao et al., 2022; Li et al., 2023a), a technique unlike Quantization-aware Training (QAT) (Liu et al., 2023; Wang et al., 2023; Xu et al., 2024) which requires heavily retraining, has been widely employed due to its efficiency and resource-friendly nature. As illustrated by Figure 3, the number of papers about PTQ for LLMs takes up nearly 70% of total quantization papers since 2022. However, despite the growing prominence of PTQ, current research still exhibits the following two limitations.
@@
Figure 1: An overview of our benchmark. To provide guidelines for future research, we first establish a comprehensive taxonomy for existing milestone PTQ methods. Then extensive and unified evaluation of the categorized PTQ strategies is provided which contains a broad range of model sizes, architectures, training levels, modalities and bitwidth. Finally, we summarize in-depth comparative analysis based on the experimental results and offer valuable recommendations for the advancement of LLM PTQ research.
@@
Firstly, the present reviews lack of in-depth insight into the characteristic of different PTQ frameworks so as to provide limited guidance for the development on advanced PTQ methods. These reviews either exhibit a lack of adequate focus on quantization (Tang et al., 2024; Yang et al., 2024), incorporate insufficient experimental setups (Gong et al., 2024; Kurtic et al., 2024), or offer inadequate analytical insights (Li et al., 2024). Admittedly, it would hold greater value for future researchers to get the information on selecting foundational PTQ frameworks for further exploration on their specific scenarios. For example, a typical question might be: Which foundational PTQ strategy should I choose to achieve better robustness across various model architecture?
Secondly, existing PTQ methods invariably prioritize enhancing the quantization performance while overlooking the trade-off among performance, model size, and quantization bitwidth. For instance, considering a intuitive questions: Which one is better, a higher-bit small model or a lower-bit large model? This is directly relevant to the selection of quantized LLMs for deployment, but previous studies cannot answer this question because they merely applied the proposed methods to models of various sizes for evaluation and presented experimental results to demonstrate their superiority without delving into the specifics.
@@
To fill these confusions, in this paper we introduce a novel PTQ benchmark for LLMs. As illustrated by Figure 1, our benchmark is built upon a comprehensive taxonomy of PTQ methods, unified evaluation including extensive experiments, and comparative analysis on the results to offer valuable recommendations. Particularly, we make full attention on weight-only PTQ methods, as they have been adopted in a greater number of practical applications, enables a wider range of bitwidth support, exhibits a richer diversity of strategies and demonstrates superior performance compared to weight-activation methods (Yuan et al., 2024). Our contributions are revealed in four dimensions:
@@
(1) Comprehensive Taxonomy. We first review the vast majority of mainstream weight-only PTQ techniques and categorize them into four classes based on their designing strategies and optimization mechanisms: compensation-based strategy represented by GPTQ (Frantar et al., 2022), optimization-based strategy exemplified by OmniQuant (Shao et al., 2023), rotation-based strategy typified by QuIP (Chee et al., 2024), and salience-based strategy characterized by AWQ (Lin et al., 2024). Our proposed taxonomy can provide a clear understanding for researchers and support our subsequent benchmarking experiments.
@@
(2) Unified Evaluation. In order to derive the characteristics of each PTQ strategy and provide reasonable and accurate recommendations, in our benchmark we conduct extensive experiments for evaluation. In detail, the experiments are mainly constructed on the most widely used open-sourced LLMs LLaMA family (LLaMA-1/2/3/3.1) with a large range of model size (7B to 70B). In order to further explore the performance, Mixture-of-Experts (MoE) LLMs, Mamba (Gu and Dao, 2023) and multimodal LLMs (MLLMs) are also included. For comprehensiveness and unification, extremely low-bit quantization (2-bit) to common 4-bit quantization are applied to all LLMs using representative baselines from the four aforementioned PTQ strategies and various evaluation tasks are covered such as generation and reasoning.
(3) Comparative Analysis. To provide guidelines for facilitating the development of more advanced PTQ methods, we summarize the characteristics of the classified four strategies based on their performance in different scenarios, including cross bitwidth/training-level robustness, cross-architecture robustness and cross modality robustness, thus offering practical recommendations for the future researchers to select foundational PTQ frameworks based on their requirements. In addition, we introduce the modelsize-bitwidth trade-off by comparing the performance of models across various sizes at different bitwidth and accordingly argue that the extremely low-bit PTQ for ultra large models need to be reexamined.
(4) Deriving from our benchmark, we further claim that a practical combination of compensation-based scheme and other PTQ strategy can achieve SOTA various robustness.
@@ 2 Taxonomy
Most previous reviews on PTQ have also categorized quantization techniques, such as symmetric or asymmetric (Gholami et al., 2022) and group-wise or channel-wise quantization (Shen et al., 2020). Such taxonomies have become increasingly inadequate to meet the research requirements of subsequent studies as PTQ techniques continue to proliferate, because such taxonomies are relatively coarse-grained, making them challenging to conduct in-depth analysis of the characteristics of each category. In this section, we compile a comprehensive list of existing mainstream weight-only PTQ algorithms and categorize them based on their underlying principles. Specifically, they are classified into four categories: compensation-based strategy, rotation-based strategy, salience-based strategy and optimization-based strategy.
@@ 2.1 Compensation-based Quantization
This strategy is pioneered by GPTQ (Frantar et al., 2022) and is one of the most influential quantization techniques currently. GPTQ first reformulates the quantization error involving the Hessian matrix and then calculates the update formula for the unquantized weights after a specific weight is quantized, which can be expressed as:
@@
where $\delta$ denotes the optimal update of the unquantized weights, $w_{q}$ is the weight as position $q$ and $\mathbf{H}$ indicates Hessian. Unlike LeCun et al. (1989) and Frantar and Alistarh (2022) which require a greedy search to identify the position that minimizes the errors for each quantization step, GPTQ partitions the weight matrix into multiple blocks and performs column-wise operations sequentially, during which the residual weights within the current block are compensated accordingly. Following GPTQ, more advanced error compensation strategies are proposed by QuantEase (Behdin et al., 2023), VPTQ (Liu et al., 2024b) and APTQ (Guan et al., 2024), achieving nearly no performance degradation even at 3-bit quantization.
@@ 2.2 Rotation-based Quantization
QuIP (Chee et al., 2024) is considered as the innovator of this strategy. Their insights reveal that quantization will be more effective when weights and proxy Hessian are incoherent, where a weight matrix $W\in\mathbb{R}^{n\times m}$ is $\mu$-incoherent if:
$\max_{i,j}|W_{ij}|=\max_{i,j}\left|e_{i}^{T}We_{j}\right|\leq\mu\|W\|_{F}/\sqrt{mn}.$
Specifically, QuIP multiplies a weight matrix by Kronecker-structured orthogonal matrices (Zhang et al., 2015) on the left and right. Such process can be thought of as a principled form of outlier reduction because the weights are similar in magnitudes, ensuring the rounding errors are not particularly large in any direction along the coordinate axes.
Inspired by QuIP, rotation-based methods rapidly gained traction. QuIP# (Tseng et al., 2024) employs Hadamard matrices (Halko et al., 2011) for rotation, achieving more efficient and superior quantization performance. QuaRot (Ashkboos et al., 2024) and SpinQuant (Liu et al., 2024c) further extend the Hadamard-rotation method to weight-activation quantization to eliminate the extreme outliers in activation channels.
@@ 2.3 Salience-based Quantization
Mixed-precision quantization methods constitute the largest subset within salience-based technique. These methods retain salient weights at higher precision while quantizing the others to lower bits. LLM.int8() (Dettmers et al., 2022) performs 8-bit quantization while preserving the weights with the top-$0.1\%$ magnitudes at 16-bit. SpQR (Dettmers et al., 2023) saves more unstructured salient weights at higher precision and employs a finer-grained group-wise quantization approach. PB-LLM (Shang et al., 2023) adopts Hessian which is considered as a more advanced metric to isolate salient weights for 8-bit quantization while binarizing the others.
Although yielding promising results, mixed-precision quantization is not hardware-friendly which may harms the inference speed and requires specialized design to accommodate the varying bit widths. Taking it into considerations, AWQ (Lin et al., 2024), the representative method of salience-based technique, is proposed. Firstly, AWQ relies on input activation as the measure instead of the previous ones that use the weights themselves as the criterion for salience. Following this, they conclude that scaling the salient weights will reduce quantization errors, thereby avoiding the deployment challenges associated with mixed-precision quantization. The scaling quantization for salient weights can be elaborated as:
@@ 2.4 Optimization-based Quantization
OmniQuant (Shao et al., 2023) is the first to introduce optimization strategy into PTQ. To avoid insufferable computational resources cost during training, an efficient block-wise learning framework is proposed, where the output of full-precision blocks serves as supervisory information to update the clipping range of the scaling factors, achieving nearly lossless performance at 4-bit quantization. The optimization objective is:
$\mathop{\arg\min}\limits_{\Theta_{1},\Theta_{2}}\|F(\mbox{W},\mbox{X})-F(Q_{w}(\mbox{W};\Theta),X)\|,$
@@
Following OmniQuant, CBQ (Ding et al., 2023) devises a two-branch framework to improve the robustness. LRQuant (Zhao et al., 2024) discovers the directional gaps between full-precision outputs and their quantized counterparts and propose a novel loss function named NLC loss to minimize the quantization error. AffineQuant (Ma et al., 2024) incorporates affine transformations into the PTQ process and optimizes the transformation matrix to reduce quantization errors.
@@
Table 1: The average perplexity($\downarrow$)/accuracy($\uparrow$) comparison results among different bitwidths, model sizes and model families of undertrained LLMs (LLaMA-1/2).
@@
Table 2: The average perplexity($\downarrow$)/accuracy($\uparrow$) comparison results among different bitwidths, model sizes and model families of fully trained LLMs (LLaMA3/3.1).
3 Benchmarking PTQ in LLMs
To clearly grasp the specific performance trait of each PTQ strategy and provide useful recommendations, in this section we conduct extensive experiments to benchmark PTQ in LLMs. The detailed experimental results are listed in each subsection with corresponding analysis, conclusions and recommendations.
3.1 Experimental Settings
In our benchmark, we select AWQ (Lin et al., 2024), GPTQ (Frantar et al., 2022), OmniQuant (Shao et al., 2023), and QuIP (Chee et al., 2024) as representatives of the four PTQ strategies, owing to their superior performance and broad practical deployment. For quantized models, we mainly focus on the LLaMA family (LLaMA-1/2/3/3.1) (Touvron et al., 2023a, b; Dubey et al., 2024), the most widely deployed open-sourced LLMs. Besides, Mixtral (Jiang et al., 2024), DeepSeeK-MoE (Dai et al., 2024), Mamba (Gu and Dao, 2023), LLaVA-1.5 (Liu et al., 2024a), and VILA-1.5 (Lin et al., 2023) are also included. The evaluation metrics contains perplexity and average zero-shot accuracy on 16 single/multi-modal reasoning tasks. Please refer to Appendix B.1 for more details.
3.2 Cross Bitwidth and Training Level Robustness
To comprehensively evaluate the four PTQ strategies, we use LLaMA families, which are the most widely deployed open-source LLMs. We present the average perplexity and accuracy in Table 1 and Table 2. For more details in each task please refer to Appendix B.2. In our exploration, we observe that even under identical experimental conditions, the performance of each PTQ strategy varies significantly across different bitwidths and training levels of LLMs.

### v4

@@ Abstract
Post-training Quantization (PTQ) technique has been extensively adopted for large language models (LLMs) compression. However, existing research lacks in-depth analysis of the strengths and applicable scenarios of different PTQ strategies, making it difficult for future researchers to choose suitable foundational framework for development based on their specific needs. To mitigate these confusions, we start by proposing a detailed taxonomy for existing mainstream methods based on their computational strategies (e.g., optimization-based, compensation-based, etc.). Subsequently, we establish a comprehensive benchmark for LLMs PTQ, named PTQ-Bench, to help select foundational framework, which evaluates the cross-bitwidth robustness, cross-structure robustness and cross-modality robustness of PTQ frameworks. Using PTQ-Bench, we conduct extensive experiments with the baseline of each categorized PTQ strategy, covering models with various sizes (7B-70B), bitwidths, structures (LLaMA1/2/3/3.1, Mixtral, DeepSeekMoE and Mamba) and modality (LLaVA1.5 and VILA1.5) on a wide range of evaluation metrics, followed by comparative analysis on the results to summarize the superior of each PTQ strategy. For example, our benchmark reveals that rotation-based technique demonstrates outstanding low-bit robustness. Finally, beyond benchmarking we present two critical implications: (1) extremely low-bit PTQ for ultra large models should be reexamined, and (2) a practical combination of compensation and other PTQ strategy can achieve SOTA various robustness. We believe our PTQ-Bench will provide valuable recommendations for future research on advanced PTQ methods. The repository for our PTQ-Bench is available at https://github.com/zjq0455/PTQ-Bench.
@@ 1 Introduction
Figure 1: An overview of our paper. To provide guidelines for future research, we first establish a comprehensive taxonomy for existing milestone PTQ methods. Then we establish a novel benchmark named PTQ-Bench for evaluating several critical characteristics of foundational PTQ strategies. Based on it, extensive and unified evaluation of the categorized PTQ strategies is provided which contains a broad range of model sizes, structures, modalities and bitwidth. Finally, we summarize in-depth comparative analysis based on the experimental results and offer valuable recommendations for the advancement of LLM PTQ research.
Large language models (LLMs) have achieved remarkable success in text generation and various reasoning tasks, with representative ChatGPT [1] and LLaMA family [65, 66, 16]. However, their massive parameter scale imposes significant memory and inference overhead, which constrain their practical deployment. To address the issue, numerous model compression techniques have been proposed, such as quantization [37, 60], pruning [19, 8, 9], low-rank decomposition [29, 75], and knowledge-distillation [23]. Among these, Post-training Quantization (PTQ) [73, 38], a technique unlike Quantization-aware Training (QAT) [46, 69, 71] which requires heavily retraining, has been widely employed due to its efficiency and resource-friendly nature. As illustrated by Figure 3, the number of papers about PTQ for LLMs takes up nearly 70% of total quantization papers since 2022.
@@
Despite the growing prominence of PTQ, current research still exhibits the following limitation: the present reviews lack of in-depth insights into the characteristic of different PTQ frameworks so as to provide limited guidance for the development on advanced PTQ methods. These reviews either exhibit a lack of adequate focus on quantization [64, 72], incorporate insufficient experimental setups [22, 34], or offer inadequate analytical insights [39]. Admittedly, it would hold greater value for future researchers to get the information on selecting foundational PTQ frameworks for further exploration on their specific scenarios. For example, a typical question might be: Which foundational PTQ framework should I choose to achieve better robustness across various model structure?
@@
To fill this confusion, in this paper we introduce a novel benchmark for foundational PTQ framework selection named PTQ-Bench. As illustrated by Figure 1, our benchmark is built upon a comprehensive taxonomy of PTQ methods, unified evaluation including extensive experiments, and comparative analysis on the results to offer valuable recommendations. Particularly, we make full attention on weight-only PTQ methods, as they have been adopted in a greater number of practical applications, enables a wider range of bitwidth support, exhibits a richer diversity of strategies and demonstrates superior performance compared to weight-activation methods [76]. Our contributions are revealed in four dimensions:
@@
(1) Comprehensive Taxonomy. We first review the vast majority of mainstream weight-only PTQ techniques and categorize them into four classes based on their designing strategies and optimization mechanisms: compensation-based strategy represented by GPTQ [20], optimization-based strategy exemplified by OmniQuant [61], rotation-based strategy typified by QuIP [7], and salience-based strategy characterized by AWQ [42]. Our proposed taxonomy can provide a clear understanding for researchers and support our subsequent benchmarking experiments.
@@
(2) PTQ-Bench for Unified Evaluation. To provide guidelines for facilitating the development of more advanced PTQ methods, we establish a comprehensive benchmark, namely PTQ-Bench, by evaluating the cross-bitwidth robustness, cross-structure robustness and cross-modality robustness of each PTQ strategy through extensive experiments to derive their characteristics. In detail, the experiments are mainly constructed on the most widely used open-sourced LLMs LLaMA family (LLaMA-1/2/3/3.1) with a large range of model size (7B to 70B). For comprehensiveness and unification, Mixture-of-Experts (MoE) LLMs, Mamba [24] and multimodal LLMs (MLLMs) are also included. Extremely low-bit quantization (2-bit) to common 4-bit quantization are applied to all LLMs using representative baselines from the each PTQ strategy and various evaluation tasks are covered such as language modeling and reasoning.
(3) Comparative Analysis. Based on the unified evaluation, we summarize the characteristics of each PTQ strategy in different scenarios, thus offering practical recommendations for the future researchers to select foundational PTQ frameworks based on their requirements. For example, we suggest to select compensation-based strategy when facing cross-structure scenario and avoid choosing naive optimization-based strategy at extremely low-bit quantization.
(4) Beyond Benchmarking. We also present the following two critical conclusion in LLMs PTQ deriving from PTQ-bench: a) Even the largest model at 2-bit performs worse than the smallest model at 4-bit in the same LLM family, so we accordingly argue that the extremely low-bit PTQ for ultra large models need to be reexamined; b) We innovatively claim that practical combination of compensation-based scheme and other PTQ strategy can achieve SOTA various robustness.
@@ 2 Taxonomy
Most previous reviews have also categorized PTQ techniques, such as symmetric or asymmetric [21] and group-wise or channel-wise quantization [62]. Such taxonomies have become increasingly inadequate to meet the research requirements of subsequent studies as PTQ techniques continue to proliferate, because such taxonomies are relatively coarse-grained, making them challenging to conduct in-depth analysis of the characteristics of each category. In this section, we compile a comprehensive list of existing mainstream weight-only PTQ algorithms and categorize them based on their underlying principles. Specifically, they are classified into four categories: compensation-based strategy, rotation-based strategy, salience-based strategy and optimization-based strategy.
@@ 2.1 Compensation-based Quantization
This strategy is pioneered by GPTQ [20] and is one of the most influential quantization techniques currently. GPTQ first reformulates the quantization error involving the Hessian matrix and then calculates the update formula for the unquantized weights after a specific weight is quantized, which can be expressed as:
@@
where $\delta$ denotes the optimal update of the unquantized weights, $w_{q}$ is the weight as position $q$ and $\mathbf{H}$ indicates Hessian. Unlike LeCun et al. [36] and Frantar and Alistarh [18] which require a greedy search to identify the position that minimizes the errors for each quantization step, GPTQ partitions the weight matrix into multiple blocks and performs column-wise operations sequentially, during which the residual weights within the current block are compensated accordingly. Following GPTQ, more advanced error compensation strategies are proposed by QuantEase [5], VPTQ [44] and APTQ [25], achieving nearly no performance degradation even at 3-bit quantization.
@@ 2.2 Rotation-based Quantization
QuIP [7] is considered as the innovator of this strategy. Their insights reveal that quantization will be more effective when weights and proxy Hessian are incoherent, where a weight matrix $W\in\mathbb{R}^{n\times m}$ is $\mu$-incoherent if:
$\max_{i,j}|W_{ij}|=\max_{i,j}\left|e_{i}^{T}We_{j}\right|\leq\mu\|W\|_{F}/\sqrt{mn}.$ (2)
Specifically, QuIP multiplies a weight matrix by Kronecker-structured orthogonal matrices [78] on the left and right. Such process can be thought of as a principled form of outlier reduction because the weights are similar in magnitudes, ensuring the rounding errors are not particularly large in any direction along the coordinate axes.
Inspired by QuIP, rotation-based methods rapidly gained traction. QuIP# [67] employs Hadamard matrices [27] for rotation, achieving more efficient and superior quantization performance. QuaRot [4] and SpinQuant [47] further extend the Hadamard-rotation method to weight-activation quantization to eliminate the extreme outliers in activation channels.
@@ 2.3 Salience-based Quantization
Mixed-precision quantization methods constitute the largest subset within salience-based technique. These methods retain salient weights at higher precision while quantizing the others to lower bits. LLM.int8() [13] performs 8-bit quantization while preserving the weights with the top-$0.1\%$ magnitudes at 16-bit. SpQR [14] saves more unstructured salient weights at higher precision and employs a finer-grained group-wise quantization approach. PB-LLM [59] adopts Hessian which is considered as a more advanced metric to isolate salient weights for 8-bit quantization while binarizing the others.
Although yielding promising results, mixed-precision quantization is not hardware-friendly which may harms the inference speed and requires specialized design to accommodate the varying bit widths. Taking it into considerations, AWQ [42], the representative method of salience-based technique, is proposed. Firstly, AWQ relies on input activation as the measure instead of the previous ones that use the weights themselves as the criterion for salience. Following this, they conclude that scaling the salient weights will reduce quantization errors, thereby avoiding the deployment challenges associated with mixed-precision quantization. The scaling quantization for salient weights can be elaborated as:
@@ 2.4 Optimization-based Quantization
OmniQuant [61] is the first to introduce optimization strategy into PTQ. To avoid insufferable computational resources cost during training, an efficient block-wise learning framework is proposed, where the output of full-precision blocks serves as supervisory information to update the clipping range of the scaling factors, achieving nearly lossless performance at 4-bit quantization. The optimization objective is:
$\mathop{\arg\min}\limits_{\Theta_{1},\Theta_{2}}\|F(\mbox{W},\mbox{X})-F(Q_{w}(\mbox{W};\Theta),X)\|,$ (4)
@@
Following OmniQuant, CBQ [15] devises a two-branch framework to improve the robustness. LRQuant [79] discovers the directional gaps between full-precision outputs and their quantized counterparts and propose a novel loss function named NLC loss to minimize the quantization error. AffineQuant [49] incorporates affine transformations into the PTQ process and optimizes the transformation matrix to reduce quantization errors.
@@
[…]
@@
3 PTQ-Bench: Benchmarking PTQ in LLMs
To clearly grasp the specific performance trait of each PTQ strategy and provide useful recommendations, in this section we first introduce our PTQ-Bench, then with its guidance we conduct extensive experiments to benchmark each PTQ strategy. The detailed experimental results are listed in each subsection with corresponding analysis, conclusions and recommendations. We also list the FP16 results of all the evaluated LLMs in Table 13 and 14 in Appendix B.3.
3.1 PTQ-Bench Setup
General setup
For assisting in selecting foundational PTQ strategies based on requirements, we select AWQ [42], GPTQ [20], OmniQuant [61], and QuIP [7] as representatives of the four PTQ strategies, owing to their superior performance and broad practical deployment. We mainly focus on the LLaMA family (LLaMA-1/2/3/3.1) [65, 66, 16], the most widely deployed open-sourced LLMs. Besides, Mixtral [32], DeepSeeK-MoE [12], Mamba [24], LLaVA-1.5 [43], and VILA-1.5 [41] are also included. The evaluation metrics contains perplexity and average inference accuracy on 16 single/multi-modal reasoning tasks. Please refer to Appendix B.1 for detailed experimental settings.
Cross-Bitwidth Robustness
Cross-bitwidth robustness quantifies an algorithm’s performance stability across varying quantization bitwidths. In PTQ-bench, we evaluate this capability by comparing the performance degradation from 4-bit to 2-bit mainly on LLaMA families.
Cross-structure Robustness
Cross-structure robustness evaluates an algorithm’s scalability and performance consistency when deployed across diverse model structures, characterizing its generalization capability beyond specific network structures. In PTQ-bench, we compare the results on MLLMs/Mamba and LLMs with traditional structures (i.e., LLaMA) to achieve this goal.
Cross-Modality Robustness
Cross-modality robustness indicates the performance consistency of an LLM PTQ method when deployed on MLLMs. In PTQ-Bench, we mainly evaluate this capability by comparing the performance trend on LLMs and MLLMs.
3.2 Cross-Bitwidth Robustness
To comprehensively evaluate the cross-bitwidth robustness of the four PTQ strategies, we use LLaMA families, which are the most widely deployed open-sourced LLMs. We present the average perplexity and accuracy in Table 1 and Table 2. In our exploration, we observe that even under identical experimental conditions, the performance of each PTQ strategy varies significantly across bitwidths.